OpenStudy (anonymous):

What is the explicit formula for the sequence? 7, 2, –3, –8, –13, . . .

OpenStudy (anonymous):

x-5 7-5=2 2-5=-3 -3-5=-8 -8-5=-13...

OpenStudy (anonymous):

That's not an option

OpenStudy (anonymous):

what are the options then

OpenStudy (anonymous):

an=-5n+12

OpenStudy (anonymous):

an=-5n+7

OpenStudy (anonymous):

an=5n-12

OpenStudy (anonymous):

an=2n-7

OpenStudy (anonymous):

an=5n-12

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yep!